Wednesday 11/25/15

Ran gels of samples B4-B10 from 2bRAD Library 1 of the PCR done on 11/24/15. Some of the PCRs either did not seem to work or might’ve washed out of the gel. I still cut out bands around 170bp for these, but will look at the quantification to see if anything is recovered. Gel slices were left in 4degC.

Samples that did not work (likely need redos): F5 8(HC1_4B), E5 (HC1_9), B5(HC2_13), B7 (HC3_9), B8(HC3_17, library 4).

 

Advertisements

Leave a Reply

Fill in your details below or click an icon to log in:

WordPress.com Logo

You are commenting using your WordPress.com account. Log Out / Change )

Twitter picture

You are commenting using your Twitter account. Log Out / Change )

Facebook photo

You are commenting using your Facebook account. Log Out / Change )

Google+ photo

You are commenting using your Google+ account. Log Out / Change )

Connecting to %s